OpenStudy (anonymous):

which number will you multiply -3/4 by equal to 1?

OpenStudy (jamesj):

Well, if you multiply by 4, you get -3, right? \[ 4 \times \frac{-3}{4} = -3 \] So if we multiply by -4 we get... \[ -4 \times \frac{-3}{4} = 3 \] So this is 3 times what we want, as we want 1. So, how to we have to modify -4 to get 1/3 of it?